# Positivstellensätze for polynomial matrices with universal quantifiers

### v2

@@ 4.3.2. Lasserre-Netzer-type Positivstellensatz for polynomial matrices with UQs
Suppose that Assumptions 3.2 and 4.1 hold, and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PSD on $\mathcal{X}\cap[-1,1]^{n}$. For any $\varepsilon>0$, there exist $\ell_{1},\ell_{2}\in\mathbb{N}$ such that for all $d\geq\ell_{1},k\geq\ell_{2}$, it holds
$F({\boldsymbol{x}})+\varepsilon\Theta_{d}I_{p}=\Sigma_{0}+\int_{\mathcal{Y}}\left\langle\Sigma,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
for some SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]_{2d}^{p}$ and $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ with $\deg_{{\boldsymbol{x}}}\Sigma\leq 2d-2d(G)$ and $\deg_{{\boldsymbol{y}}}\Sigma\leq 2k$.
Proof.
By Proposition 4.2, there exist $\ell_{1},\ell_{2}\in\mathbb{N}$ such that for all $d\geq\ell_{1}$ and $k\geq\ell_{2}$, there is a feasible point $(r,\Sigma_{0}^{\prime},\Sigma)$ of (6) with $r\leq\varepsilon$. We have
$F({\boldsymbol{x}})+r\Theta_{d}I_{p}=\Sigma_{0}^{\prime}+\int_{\mathcal{Y}}\left\langle\Sigma,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
with $\deg(\Sigma_{0}^{\prime})\leq 2d$, $\deg_{{\boldsymbol{x}}}\Sigma\leq 2d-2d(G)$, and $\deg_{{\boldsymbol{y}}}\Sigma\leq 2k$. Then, for all $d\geq\ell_{1}$ and $k\geq\ell_{2}$, it holds
$F({\boldsymbol{x}})+\varepsilon\Theta_{d}I_{p}=F({\boldsymbol{x}})+r\Theta_{d}I_{p}+(\varepsilon-r)\Theta_{d}I_{p}=(\Sigma_{0}^{\prime}+(\varepsilon-r)\Theta_{d}I_{p})+\int_{\mathcal{Y}}\left\langle\Sigma,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
The conclusion then follows. ∎
In case that UQs are not present, we obtain the following corollary.
Corollary 4.4.
Let $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{q}$ and $\mathcal{W}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H({\boldsymbol{x}})\succeq 0\}.$ Suppose that there are positive number $\lambda>0$, and open and bounded subset $\mathcal{O}$ of $\mathcal{X}$ such that $H({\boldsymbol{x}})\succeq\lambda I_{q}$ on $\mathcal{O}$. If $F({\boldsymbol{x}})$ is PSD on $\mathcal{W}\cap[-1,1]^{n}$, then for any $\varepsilon>0$, there exists some $\ell\in\mathbb{N}$ such that for all $d\geq\ell$, it holds
$F({\boldsymbol{x}})+\varepsilon\Theta_{d}I_{p}=\Sigma_{0}+\left\langle\Sigma,G\right\rangle_{p}$
for some SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]_{2d}^{p}$ and $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}}]_{2d-2d(G)}^{pq}$.
5. Applications to robust PMI constrained optimization
Verifying PMIs over a prescribed set has a wide range of applications in many fields. For instance, many control problems for systems of ordinary differential equations can be formulated as convex optimization problems with PMI constraints that must be satisfied over a specified portion of the state space [15, 16, 19, 40, 51]. These problems are typically formulated as follows:
$\inf_{{\boldsymbol{\gamma}}\in{\mathbb{R}}^{r}}\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\ \text{ s.t. }\ P({\boldsymbol{x}},{\boldsymbol{\gamma}})\coloneqq P_{0}({\boldsymbol{x}})-\sum_{i=1}^{r}P_{i}({\boldsymbol{x}})\gamma_{i}\succeq 0,\ \forall{\boldsymbol{x}}\in\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H({\boldsymbol{x}})\succeq 0\},$
where ${\boldsymbol{c}}\in{\mathbb{R}}^{r}$, $P_{0},\ldots,P_{r}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $H\in{\mathbb{S}}[{\boldsymbol{x}}]^{q}$.
However, due to estimation errors or lack of information, the data of real-world problems often involve uncertainty. Therefore, ensuring the robustness of PMIs over a prescribed set under uncertainty is a critical issue, which could be formulated as the following robust optimization problem:
$\tau^{\star}\coloneqq\inf_{{\boldsymbol{\gamma}}\in{\mathbb{R}}^{r}}\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\ \text{ s.t. }\ P({\boldsymbol{x}},{\boldsymbol{\gamma}})\succeq 0,\ \forall{\boldsymbol{x}}\in\mathcal{X}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid G({\boldsymbol{x}},{\boldsymbol{y}})\succeq 0,\ \forall{\boldsymbol{y}}\in\mathcal{Y}\},$ (8)
Consequently, Positivstellensätze for polynomial matrices with UQs developed in this paper serve as a powerful mathematical tool for addressing this issue. Indeed, by leveraging the SOS-structured certificates provided by Positivstellensätze for positive definiteness of $P({\boldsymbol{x}},{\boldsymbol{\gamma}})$ over $\mathcal{X}$, we are able to establish converging hierarchies of SDP relaxations for the problem (8).
Concretely, if $\mathcal{X}$ is compact, then one could apply the Scherer-Hol-type Positivestellensatz (Theorem 3.1) to construct a hierarchy of SDP relaxations for (8):
$\tau_{k}\coloneqq\inf_{{\boldsymbol{\gamma}},\Sigma_{0},\Sigma}&\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\\ \text{s.t.}&\ P({\boldsymbol{x}},{\boldsymbol{\gamma}})=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),\\ &\ \Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\,\,\text{are SOS matrices},\\ &\ \deg\Sigma_{0},\ \deg_{{\boldsymbol{y}}}\Sigma\leq 2k,\ \deg_{{\boldsymbol{x}}}\Sigma\leq 2k-\deg_{{\boldsymbol{x}}}G.\end{aligned}\right.$
It is clear that the sequence $(\tau_{k})_{k\in\mathbb{N}}$ is non-increasing upper bounds of $\tau^{\star}$. Assuming that the Slater condition holds for (8), i.e., there exists a point $\bar{{\boldsymbol{\gamma}}}\in{\mathbb{R}}^{r}$ such that $P({\boldsymbol{x}},\bar{{\boldsymbol{\gamma}}})\succ 0$ for all ${\boldsymbol{x}}\in\mathcal{X}$, the convergence of $\tau_{k}\to\tau^{\star}$ as $k\to\infty$ is guaranteed under Assumptions 2.1 and 3.2 by Theorem 3.1.
Other Positivstellensätze presented in this paper can also be employed to derive the corresponding hierarchies of SDP relaxations for (8). The study of convergence of these hierarchies provides an intriguing avenue for future research. For instance, if $\mathcal{X}$ is non-compact, then one could apply the Putinar–Vasilescu-type Positivstellensatz for the inhomogeneous case (Corollary 4.1) with a fixed $\varepsilon>0$ to construct the following hierarchy of SDP relaxations for (8) ($d_{P}\coloneqq\max_{i,j\in[p]}\lfloor\deg_{{\boldsymbol{x}}}(P_{ij})/2\rfloor+1$):
$\tau_{k}(\varepsilon)\coloneqq\inf_{{\boldsymbol{\gamma}},\Sigma_{0},\Sigma}&\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\\ \text{s.t.}&\ \theta^{k}(P({\boldsymbol{x}},{\boldsymbol{\gamma}})+\varepsilon\theta^{d_{P}}I_{p})=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),\\ &\ \Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\,\,\text{are SOS matrices},\\ &\ \deg\Sigma_{0},\ \deg_{{\boldsymbol{y}}}\Sigma\leq 2(k+d_{P}),\ \deg_{{\boldsymbol{x}}}\Sigma\leq 2(k+d_{P}-d_{G}).\end{aligned}\right.$
Assuming that $\tau^{\star}$ is attainable and Assumption 3.2 holds, by Corollary 4.1, there exists $\bar{k}\in\mathbb{N}$ such that $\tau_{k}(\varepsilon)\leq\tau^{\star}$ for all $k\geq\bar{k}$. How to bound the number $\bar{k}$ from above, and whether $\tau_{k}(\varepsilon)\to\tau^{\star}$ as $\varepsilon\to 0$ and $k\to\infty$ under certain conditions, could be interesting topics for further study.

### v4

@@ 5.3. The non-compact case
Let the Archimedean condition hold for the quadratic module $\mathcal{Q}^{pq}(\mathbf{H})$. Then the limit $\tilde{\tau}_{0}(\infty,\infty)$ exists and $\tilde{\tau}_{0}(\infty,\infty)\leq\tau^{\star}$. Moreover, if the optimal solution set $\mathcal{S}$ of (15) is nonempty and bounded, we have $\tilde{\tau}_{0}(\infty,\infty)=\tau^{\star}$.
Note that we set the same degree bounds on $\boldsymbol{x}$} and $\boldsymbol{y}$} in the representations of $\Lambda({\boldsymbol{x}},{\boldsymbol{y}})$ in (13) and (16), so that we can compare the behaviors of $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ and $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ when applying both (13) and (16) to (8) if $\mathcal{X}=\mathcal{X}\cap{\mathbb{R}}^{n}_{+}$.
Example 5.3.
Consider the problem (8) where we let $m=1$, $n=r=2$, $\mathcal{Y}=[0,1]$,
$G({\boldsymbol{x}},y)=Q_{1}^{\intercal}\left[\begin{array}[]{cc}x_{1}+x_{2}&0\\ 0&-2y(1-y)+(1-y)^{2}x_{1}+y^{2}x_{2}\end{array}\right]Q_{1},$
and
$\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})=Q_{2}^{\intercal}\left[\begin{array}[]{cc}-2+\left(-\frac{\gamma_{1}}{2}+\frac{3\gamma_{2}}{2}+3\right)x_{2}+\left(\frac{3\gamma_{1}}{2}-\frac{\gamma_{2}}{2}+3\right)x_{1}&0\\ 0&-2-\gamma_{1}x_{2}-\gamma_{2}x_{1}\end{array}\right]Q_{2}.$
Here, $Q_{1}$ and $Q_{2}$ are two fixed arbitrary $2\times 2$ orthogonal matrices.
For any fixed $(x_{1},x_{2})$, consider the function
$f(y)=-2y(1-y)+(1-y)^{2}x_{1}+y^{2}x_{2}=(x_{1}+x_{2}+2)y^{2}-2(x_{1}+1)y+x_{1}.$
The evaluations $f(0)=x_{1}$ and $f(1)=x_{2}$ imply that $x_{1}\geq 0$ and $x_{2}\geq 0$. Assuming $x_{1},x_{2}\geq 0$, the minimum of $f(y)$ is
$f\left(\frac{x_{1}+1}{x_{1}+x_{2}+2}\right)=\frac{x_{1}x_{2}-1}{x_{1}+x_{2}+2}\ \text{ where }\ \frac{x_{1}+1}{x_{1}+x_{2}+2}\in\mathcal{Y}.$
Thus we have
$\mathcal{X}=\{(x_{1},x_{2})\in{\mathbb{R}}^{2}\mid x_{1}\geq 0,\ x_{2}\geq 0,\ x_{1}x_{2}\geq 1\}.$
Then similarly, we can show that $\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})\succeq 0$ for all ${\boldsymbol{x}}\in\mathcal{X}\cap{\mathbb{R}}_{+}^{n}$ if and only if
$2+\left(-\frac{\gamma_{1}}{2}+\frac{3\gamma_{2}}{2}+3\right)x_{2}+\left(\frac{3\gamma_{1}}{2}-\frac{\gamma_{2}}{2}+3\right)x_{1}\geq 0\ \text{ and }\ -2-\gamma_{1}x_{2}-\gamma_{2}x_{1}\geq 0$
on $\mathcal{X}$ if and only if
$\frac{\gamma_{1}}{2}-\frac{3\gamma_{2}}{2}-3\leq 0,\ -\frac{3\gamma_{1}}{2}+\frac{\gamma_{2}}{2}-3\leq 0,\ \phi_{1}({\boldsymbol{\gamma}})\coloneqq\left(\frac{\gamma_{1}}{2}-\frac{3\gamma_{2}}{2}-3\right)\left(-\frac{3\gamma_{1}}{2}+\frac{\gamma_{2}}{2}-3\right)-1\geq 0,$
and
$\gamma_{1}\leq 0,\ \gamma_{2}\leq 0,\ \phi_{2}({\boldsymbol{\gamma}})\coloneqq\gamma_{1}\gamma_{2}-1\geq 0.$
The feasible set $\mathcal{F}$ is shown in light blue in Figure 2. Since $\mathcal{X}=\mathcal{X}\cap{\mathbb{R}}^{2}_{+}$, we can apply both SDP relaxations (13) and (16) to this problem and compare the behaviors of $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ and $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$.
Consider the objective function ${\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}=\gamma_{1}-\gamma_{2}$. One can check that the optimal solution is one of the solutions of $\phi_{1}({\boldsymbol{\gamma}})=\phi_{2}({\boldsymbol{\gamma}})=0$:
${\boldsymbol{\gamma}}^{\star}=\left(1-\sqrt{5}-\sqrt{5-2\sqrt{5}},1-\sqrt{5}+\sqrt{5-2\sqrt{5}}\right)\approx(-1.9626,-0.5095),$
and the optimal value $\tau^{\star}\approx-1.4531$. The computational results for $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ and $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ with different $\varepsilon$ and $(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ are presented in Table 4. As expected, for a given $\varepsilon>0$, both $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ and $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ provide lower bounds for the optimal value $\tau^{\star}$ when $k_{{\boldsymbol{x}}}$ and $k_{{\boldsymbol{y}}}$ are sufficiently large (except for $\tau_{0.01}(2,15)$ and $\tilde{\tau}_{0.01}(2,15)$ which requires larger $k_{{\boldsymbol{x}}}$ and $k_{{\boldsymbol{y}}}$ as $\varepsilon=0.01$ is small). Moreover, for a fixed $\varepsilon>0$, both quantities are non-increasing as $k_{{\boldsymbol{x}}}$ and $k_{{\boldsymbol{y}}}$ increase, and for fixed $k_{{\boldsymbol{x}}}$ and $k_{{\boldsymbol{y}}}$, they are non-decreasing as $\varepsilon$ decreases. Compared with $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$, the values $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ with the same $(\varepsilon,k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ are smaller and can be computed in much less time.
To illustrate the behavior of optimal solutions of (13) and (16) under varying $\varepsilon$ and $(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$, like in Example 5.1, we solve (13) and (16) with the objective function $\gamma_{1}\cos\theta+\gamma_{2}\sin\theta$ for $40$ evenly spaced values of $\theta$ over $0,2\pi]$, and extract the corresponding optimal solutions. The feasible sets of (13) and (16), which serve as approximations of $\mathcal{F}$, are then outlined by connecting these optimal solutions and shown in Figure 2. As one can see, the behavior of these optimal points aligns with the quantities $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ and $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$, confirming the theoretical expectations discussed above. In particular, it is evident from the figure that the feasible set of (16) contains that of (13) with the same $(\varepsilon,k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$, which corresponds to the fact that $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ is smaller than $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$. ∎
Table 4. Computational results for Example 5.3.
$\varepsilon$ $(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}}$)=(2, 5) $(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}}$)=(2, 10) $(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}}$)=(2, 15)
$\tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tilde{\tau}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tilde{\tau}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tilde{\tau}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time
$0.1$ -1.3716/2.20s -1.4922/1.69s -1.4844/32.3s -1.6587/7.32s -1.5650/547s -1.6837/27.2s
$0.05$ -1.2664/2.34s -1.3163/1.93s -1.3698/31.1s -1.4899/7.21s -1.4552/507s -1.5230/29.9s
$0.01$ -1.1873/2.57s -1.1806/2.17s -1.2949/33.3s -1.3625/7.86s -1.3721/451s -1.3900/35.7s
Figure 2. The feasible set (in light blue) of the problem in Example 5.3 and the feasible sets of its SDP relaxations (13) (outlined with blue line) and (16) (outlined with red line) with the same values $(\varepsilon,k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$.